OpenStudy (anonymous):

Implicitly differentiate y+sqrt(y)*ln(x)=x^2+y^2 and simplify. Show all steps.

OpenStudy (anonymous):

I have (dy/dx)+1/2sqrt(y)*(dy/dx)*ln(x)+sqrt(y)/x=2x+2y(dy/dx) need help finding dy/dx

OpenStudy (anonymous):

Is that:\[y+\sqrt y *\ln(x)=x^2+y^2\]

OpenStudy (anonymous):

yes